# Flips in colorful triangulations

### v2

@@ 8. Open questions
Tables 3 and 3 show all coloring sequences $\alpha$ on up to $N\leq 11$ points for which the graph ${\mathcal{F}}_{\alpha}$ has no Hamilton path or cycle. The sequences are shown up to rotational symmetry, reversal, and exchange of the two colors. In several cases, Theorems 3, 4 and 5 provide an explanation for the non-Hamiltonicity; see the third column in the tables. In the other cases, we are still missing such an explanation.
@@
Another interesting question concerns bijections between $k$-ary trees and classes of permutations. For $k=2$ (binary trees), there is a natural bijection to 231-avoiding permutations. Are there similar correspondences between $k$-ary trees and pattern-avoiding permutations for $k\geq 3$? In particular, do tree rotations translate to nice operations on the permutations, specifically to so-called jumps heavily used in [HHMW22, HM21, MM23, CMM24]?

### v3

@@ 8. Open questions
Tables 4 and 4 show all coloring sequences $\alpha$ on up to $N\leq 11$ points for which the graph ${\mathcal{F}}_{\alpha}$ has no Hamilton path or cycle. The sequences are shown up to rotational symmetry, reversal, and exchange of the two colors. In several cases, Theorems 3, 4 and 5 provide an explanation for the non-Hamiltonicity; see the third column in the tables. In the other cases, we are still missing such an explanation.
@@
Another interesting question concerns bijections between $k$-ary trees and classes of permutations. For $k=2$ (binary trees), there is a natural bijection to 231-avoiding permutations. Are there similar correspondences between $k$-ary trees and pattern-avoiding permutations for $k\geq 3$? In particular, do tree rotations translate to nice operations on the permutations, specifically to so-called jumps heavily used in [HHMW22, HM21, MM23, CMM25]?